Tuesday, 1 August 2017


DAY 20: 24/07/17
The results from DAY 19's PCR is shown in figure 16.

It can be seen from figure 16 that the PCR with primers NRP1-K185-REV didn't work on DNAs #5-1, #5-2, #6-3, #6-4, even ad adapted conditions (lower annealing temperature, 48oC). So, a new PCR was prepared with NRP1-HBT-REV primer mix instead. Results are shown in figure 17.


It can be seen from figure 17 that there is a very light DNA band being emitted for #6-3 and #6-4. However, it is not very clear, so a new set of results were ran with the resting PCR. This time the exposure was changed in order to try and see the DNA bonds more clearly. Results are shown in figure 18.


It can be seen from figure 18 that there are DNA bands, however, the signal is weak (hence the need to increase exposure). Figure 19 is the same results as shown in figure 18, but with lowered exposure, showing no DNA bands at all to the naked eye (I thought this was a cool thing to show.. how looking at things from different 'angles' can change your interpretation of things).


From figure 15, it can be seen that N5-3 and N5-4 give a nice DNA signal, so those 2 PCR were sent to be sequenced against HBS2 primer.

DAY 21: 25/07/17
Today I will practice to do serial dilutions and spotting assays using a 96 well plate and a pinning tool (apparently this tool is very expensive, so I shall have to be extra careful with it)

Dr. MacNeill prepared beforehand 3 yeast cultures: wild type, N5-1 and N5-3. In order to do the dilutions, the OD600nm of each culture was measured, giving wild type OD = 0.26, N5-1 = 0.24, and N5-3 OD = 0.23. As it can be seen, the OD values for all 3 yeast cultures are of a close range, so there was no need to dilute the cells to OD600nm of 0.1 (also, this is just a practice day). Once the OD values were figured out, the cell dilutions was prepared, and cells were pinned to YES4 plates using the pinning tool (exciting bit). Dr MacNeill showed me step by step how do all of this, because I will have to do it on my own from no on.

In preparation for tomorrow, I prepared a culture for N5-1 and wild type (WT) pre-cultures (previously prepared by Dr. MacNeill), and left it to grow overnight at a shaking bath at 30oC.

DAY 22: 26/07/17
Today I did everything I learnt yesterday (on my own) using the N5-1 and WT cultures I prepared yesterday. The plates were incubated at 20oC, 32oC and 35oC to see how temperature influences growth.

The OD600nm values for the WT and N5-1 cultures were 0.17 and 0.19, respectively. Therefore, there wasn't a need to dilute the solutions to an OD value of 0.1.

Today pre-cultures for WT, N1, N3, N5, N8 and N10 DNAs were prepared and allowed to grow overnight at 32oC.

DAY 23: 27/07/17
Today I made cultures from the pre-cultures prepared yesterday, and the cultures were allowed to born overnight at a 30oC bath with shaking.

The plates prepared on DAY 21 and 22 were analysed. They still need to be left to grow for longer, however, it was possible to see the effects of the dilutions and how well presented a plate is by using a pinning tool (instead of manually pipetting all solutions into the agar plate). See this photo so you know what I mean:


DAY 24: 28/07/17
So today I used the WT, N1, N3, N5, N8, and N10 yeast cultures, prepared yesterday, to do serial dilutions and spotting assays.

The OD600nm values for the WT, N3 and N8 yeast cultures were below 0.1, so they were allowed to grow for ~30/40 minutes longer in the shaking water bath at 30oC. Apart from culture N10, all the other cultures, N1, N5 had their OD values of 0.11 and 0.13 (slightly above from 0.1, the desired OD value). All cultures were adjusted to have an OD value of 0.1, either by diluting or letting the cells grow more. Once all cultures had their OD value of 0.1, the serial dilutions were prepared

The serial dilution solutions were pinned on YE4S plates + HU (hydroxyurea) plates, with different HU concentrations: 0, 3, 6, 9 and 12 mM. Solutions were also pinned on YE4S plates to undergo UV light treatment. The UV light wavelengths were 50, 100, 150, 200 and 250nm. All plates were incubated at 32oC.

INFO: The purpose of growing cell at 'damaging' conditions (hydroxyure, UV light) is to see if the mutated cells (N1, N3, N5, N8, N10) are able to grow and if so, when they stop being able to grow.


DAY 16: 17/07/17
After a 2 weeks break, where I went to London to visit my friend, I'm back in the lab, and excited for the next part of my internship project.

Today I did miniprep DNA for colony PCR using N1, N3, N5, N8 and HBT plasmids.

INFO: what was previously called plasmids 1, 3, 5, 8, 10 and b3.4 are now called N1, N3, N5, N8, N10 and HBT, respectively.

The PCR was prepared in the following order: MW maker, N1, N3, N8, N10, N5 and HBT, where NSEQ1-RS1 primer mix was used for N1, N3, N8 and N10 DNAs; FS1-RS2 primer mix used for N5 DNA; and 5NSEQ-HBT-PCR-REV primer mix used for HBT DNA.

DAY 17: 18/07/17
The PCR prepared yesterday was ran on agarose gel to see the results. The results are shown in figure 10.
Figure 10: Agarose gel analysis of DNA MW marker against PCR solutions N1, N3, N8, N10, N5 and HBT, respectively

As it can be seen from figure 10, the PCR didn't really work for most of the DNA samples, since only the primers are giving out a signal for DNA samples N1, N3, N8, N10. However, for DNA sample N5, there are two visible strands, suggesting that PCR has fully worked, showing the prier strand and the DNA part of interest strand.

To check whether the DNA samples were ok, a new PCR was prepared. This time, using FS1-RS2 and NSEQ1-RS1 primer mixes. Results are shown below in figure 11.

Figure 11: Agarose gel analysis of DNA MW marker against PCR solutions N1, N3, N8, N10, N5, HBT (treated with FS1-RS2 primer mix), N1, N3, N8, N10, N5, HBT (treated with NSEQ1-RS1 primer mix), respectively.

As it can be seen from figure 11, the DNA samples are ok because, at least for FS1-RS2 primers, all DNA samples gave signal at the point of interest, suggesting the primers were the problem on the previous PCR reaction. Furthermore, it can be seen that the solutions treated with NSEQ1-RS1 primer mix didn't give signal for the DNA area of interest.

To check whether it was the primer not working, a new primer mix with NSEQ1-RS1 primers was prepared to do a new PCR.

DAY 18: 19/07/17
The results from yesterday's PCR is shown in figure 12.


It can be seen from figure 12 that the PCR didn't work, as only the primers themselves are emitting signal, thus, suggesting that the NSEQ1-RS1 primers combination is not binding properly to the DNA's part of interest.

In order to identify which primer is not working properly, new primer combinations were prepared and tested on N1 and N3 DNAs. The first new primer mix combination is NSEQ1-RS2 (RS2 have worked before with FS1 primer, as shown in figure 10 and 11). The second new primer mix combination is 5NSEQ-RS1. Results are shown in figure 13.


From figure 13, it can be seen that N1 DNA treated with NSEQ1-RS2 primers, show a very light DNA band. This suggests that some DNA from the part of interest is giving out signals. Furthermore, N3 DNA treated with NSEQ1-RS2 primers did not give any signal other than the primer itself. Neither N1 or N3 DNAs, treated with 5NSEQ-RS1 primers, gave any signals. These results suggest that the NSEQ1 primer is the problem in the NSEQ1-RS2 primer mix and we don't have enough data to determine which primer is the problem from the 5NSEQ-RS1 primer mix (results from figure 10, regarding the HBT treated with 5NSEQ/HBT-PCR-REV were inconclusive).

In other to obtain more data, other colonies of the HBT  tag colonies were treated with 5NSEQ-HBT0PCR-REV primer mix. Results are shown in figure 14.

INFO: The primer mix should only bind if the HBT tag is present in the DNA.

Figure 14: Agarose gel analysis of DNA MW marker against HBT tag colonies 3.1, 3.2, 3.4, 3.5, 3.6, treated with 5NSEQ-HBT-PCR-REV primer mix.

It can be seen from figure 14 that out of 5 HBT tag colonies, only one seems to actually have the HBT tag. The colony with a HBT tag seems to be colony #5-2.

DAY 19: 20/07/17
The only colony with a HBT tag seems to be colony #5-2, in figure 14, but since none of the others appear to have the HBT tag, which they should have (in theory), a new PCR with all #5 colonies and #6 colonies, as a control, was ran with different primers. The colonies were treated with NRP1-K185-REV primer mix, and also NRP3-NRP4 primer mix. There are two colonies for #5 and two for #6, so they were labelled #5-1, #5-2, #6-3, #6-4, and each colony was treated with each set of primer mix.

Furthermore, from figure 10, it can be seen that only N5 had a DNA signal, so DNA from all N5 and N6 colonies were prepared and treated with FS1-RS2 primer mix for PCR. There are four colonies within N5 and four within N6, so they were labeled N5-1, N5-2, N5-3, N5-4, N5-6, N6-6, N6-7 and N6-8.

The results for all #5, #6, N5 and N6 colonies are shown in figure 15.


From figure 15 results, it can be seen that the NRP1-K185-REV primer mix didn't work, as there is no DNA signal visible. However, that was expected because primer K185-REV optimal annealing temperature is lower than 60oC, so a new PCR with annealing temperature at 48oC was ran for #5-1, #5-2, #6-3 and #6-4 DNAs treated with NRP1-L185-REV.


DAY 13: 26/06/17
Today I will do colony PCR using 1, 5, 8 and 10 DNAs. This is to ensure DNA is going out signals. The gel was wrongly ran at 20V instead of 100V, but that's okay because the whole purpose of this PCR was to ensure that the DNA present is strong enough to give out a signal, and it is, as we can see in the image below.

DAY 14: 27/06/17
Today the only thing I had to do was to plate out the b3.4 1/100 dilution and 3 1/100 dilution grown solutions.

Plate out the solutions were fairly quick, so to do more things in the lab, I helped Dr. MacNeill with his own research buy doing colony PCR with some of his plasmids DNAs.

DAY 15: 28/06/17
Today PCR was ran using 1, 5, 8 and 10 plasmid DNAs with NRP1-NRP2 and NRP3-NRP4 mix primers. Results are shown in figure 7.


As we can see from figure 7, PCR results suggest that the PCR didn't work, so everything was repeated with optimising conditions, until it worked and the DNA gave out signals.

➢ A 'wildtype' was induced and a new primer replaced NRP2 was used: K185-REV primer (figure 8)
➢ Optimising conditions also involved changing the annealing temperature in the PCR reaction (annealing temperature went from 60oC to 48oC) (figure 9)

Results are shown in figure 8 and 9.


Ok, so Dr MacNeill is going on holidays with his wife, which means, I get some free time myself!!! 😚😃


DAY 20: 24/07/17 The results from DAY 19's PCR is shown in figure 16.                                       It can be seen from figu...